OpenStudy (anonymous):

What is the equation of the line which includes points (2, 3) and (0, 10)?

OpenStudy (anonymous):

find the slope, then write an equation in slope-intercept. plug in on of the coordinates, and you will have the equation

OpenStudy (anonymous):

slope: (10-3)/(0-2) -7/2

OpenStudy (anonymous):

Slope = -7/2

OpenStudy (gorv):

general equation of line y=mx+c put both points (2,3) 3=2m+c c=3-2m put (0,10) 10=0+c c=10 10=c=3-2m 10-3=-2m m=-7/2 so equation is y=-7x/2+10

OpenStudy (anonymous):

since we know that the y-int is 10 the equation will be y=-7x/2 + 10

OpenStudy (anonymous):

The equation is f(x)= -7/2x + 10

OpenStudy (anonymous):

@gorv You didn't need to plug in 0 and 10, because the y-int(b) is just where x = 0

OpenStudy (gorv):

i knw...i was just telling him @greenlegodude57